OpenStudy (anonymous):

Find the distance between the points (4, 2) and (-8, 7). A. 13 B. 15 C. 12 D. 11

OpenStudy (anonymous):

distance formula is \[\sqrt{(x2-x1)^{2}+(y2-y1)^{2}}\]

OpenStudy (nurali):

d=sqrt((x2-x1)^2+(y2-y1)^2) (x1,y1)=(4, 2) (x2,y2)= (-8, 7) d=sqrt((-8-4)^2+(7-2)^2) d=sqrt((-12)^2+(5)^2) d=sqrt(144+25) d=sqrt(169) d=13